Name the following metal complexes or complex ions: formula name [ZnF] (en)2] [ TiBrz(co)] [c: (OH)(NH3), )* x 6 ?
Answer #1

The name of the given coordination complexes can be determined by using some basic set of rules as discussed in the pic attached!coordination compounds a some important rules for naming of used here are as follows: Ц (1) For naming of the coordination sp2 [Ti Bra (con Ligands are named in order : Bromo, carbonye. o.s of Ti :- +2 24-3=-1 in=-1+3 I =+2 Name : Tribromo caubonyeti